# Josephson anomalous vortices

### v1

@@ Abstract
with $\mu_{0}$ the vacuum permeability, and $R$ the radius of the vortex. Substituting in $R=\xi_{0}$, we find a maximum magnetic field strength of $B_{\rm anom}=2\pi(R/\xi_{0})(\alpha\xi_{0})^{2}(\eta/\Delta_{0}\xi_{0}^{4})B_{0}$, some 1.5 times larger than what is seen in numerics.
For very small ($L_{x}\times L_{y}=0.1\xi_{0}\times 0.1\xi_{0}$) junctions the system is dominated by the antivortex phase, and so the magnetic field at the center is usually small. For large $\alpha$ the system is close to the transition and so due to finite size we find a small magnetic field. For large junctions ($L_{x}\times L_{y}=5\xi_{0}\times 5\xi_{0}$) the parameter space is dominated by the vortex phase and the transitional phase. Large counterpropagating currents around the terminals counteract the magnetic field due to the central vortex (see Sec. S3 in [22]). For $L_{x}\times L_{y}=0.5\xi_{0}\times 0.5\xi_{0}$ junctions the parameter space is evenly mixed between the antivortex, vortex, and transitional phases. At around $J=\Delta_{0}$ there is a large magnetic field due to this transition (see Sec. S3 in [22]). When the system is only a little larger ($L_{x}\times L_{y}=\xi_{0}\times\xi_{0}$), the parameter space is dominated by the vortex phase, with a small transitional phase. At around $J=\Delta_{0}$ there is a large magnetic field of around $B_{\rm anom}=0.04B_{0}$ due to the vortices, in agreement with the simple approximation above. Because $f_{z}\sim J/(J^{2}+\omega^{2})$ (with $\omega$ the Matsubara frequency) [22], we expect the magnetic field strength to peak at some finite $J$, and decay for large $J$; in the numerics, this peak occurs around $J\approx\alpha/2$.
@@
Characterization and manipulation— Because circulating currents produce a magnetic field, a natural approach to characterizing these vortices is via nano-magnetometry. Scanning probe techniques are well-established and can operate at cryogenic temperatures with high spatial resolution. Magnetic force microscopy has already been used to study conventional Josephson vortices [27, 28], and can operate with a resolution of 10 — 50 nm [29], or even atomic resolution [30]. Scanning nitrogen-vacancy center [31] and nanoSQUID methods [32] can now both achieve spatial resolutions of around 50 nm. These magnetometry techniques are complicated by the need to subtract the background ferromagnetic signal. One possible solution is to include a back gate to control the Rashba interaction [33]. One could then effectively switch off SOC to isolate the background ferromagnetic signal. The absence of SOC can, for example, be confirmed via the absence of a Josephson diode effect.
Such gate-based techniques for controlling Rashba SOC could also be used to manipulate these vortices. Consider a large weak link which has only weak (or absent) Rashba SOC. One can define an island within the weak link with strong Rashba SOC via a small back gate. We show in Fig. missing 3 examples where such Rashba “islands” can be used to control the size, position, and shape of a JAV. Such methods could be used to realize JAVs in arbitrarily sized junctions, overcoming the necessity of defining a square geometry of a coherence length. Indeed, several different vortices could be defined within a single weak link using this technique.
Superconducting textures— All components of the superconducting correlations must respect the 180°$\degree$ rotation symmetry of the junction. This requirement, in combination with the junction geometry and spin-orbit coupling, allows for phase textures of $f_{z}$, such as shown in Fig. missing 4(a). This texture is characterized by phase gradients in both the $x$ and $y$ directions. We find similar textures for all system sizes studied, above a threshold $J,\alpha$. We also point out that the $f_{x}$ and $f_{y}$ components can be finite in superconducting heterostructures with both SOC and ferromagnetism.
Consider a ferromagnetic disk with strong SOC and a superconducting rim. Due to the rotational symmetry $f_{s}$ and $f_{z}$ can only have radial dependence. In contrast, $f_{x}$ and $f_{y}$ can have both radial and angular dependence, implying a texture could form. We solve the full problem numerically for such a disc (Fig. missing 4(b)) and also for a square geometry (Fig. missing 4(c)), such as the one studied in Figs. 1-3. Indeed, we find a striking resemblance between these $f_{x}-f_{y}$ textures and magnetic textures. It is currently unclear how to observe these vortices.
Figure 4: (a) Phase texture in a square geometry, computed numerically. At $x\approx 0$, $f_{z}=0$. Superconducting $f_{x},f_{y}$-textures computed (b) numerically on a square geometry ($L_{x}=L_{y}=\xi_{0}$), and (c) numerically on a disc geometry, with the rim superconducting and the bulk of the disc a ferromagnet with SOC (radius $R=\xi_{0}$). We plot the components at the smallest Matsubara frequency. Parameters: $(J,\alpha)=(\Delta_{0},2/\xi_{0})$.
Outlook— We describe the characterization and manipulation of Josephson anomalous vortices, which appear in Josephson junctions where the weak link is a weak ferromagnet with strong spin-orbit coupling. The exchange field and spin-orbit coupling conspire to lead to circulating currents that we associate phenomenologically to a previously unexplored rotary term in the free energy. We also find that the same ingredients lead to textures in the triplet components ${\bm{f}}_{t}$, which do not directly contribute to the current.

### v2

@@ Abstract
with $\mu_{0}$ the vacuum permeability, and $R$ the radius of the vortex. Substituting in $R=\xi_{0}$, we find a maximum magnetic field strength of $B_{\rm anom}=2\pi(R/\xi_{0})(\alpha\xi_{0})^{2}(\eta/\Delta_{0}\xi_{0}^{4})B_{0}$, some 1.5 times larger than what is seen in numerics for vortices in the $(1\times 1)\xi_{0}^{2}$ system.
For the $(1\times 1)\xi_{0}^{2}$ system, which hosts vortices for most of parameter space, we see the largest ${\bm{B}}_{\rm anom}$, increasing with increasing $\alpha$, and peaking at $J\sim\Delta_{0}$. This magnetic field is due to the vortices. For the $(0.5\times 0.5)\xi_{0}^{2}$ system we see a weaker magnetic field, also increasing with $\alpha$ and peaked at $J\sim\Delta_{0}$. This magnetic field is due to the mixture of vortex and antivortex patterns around the phase transition. Surprisingly, for the $(0.1\times 0.1)\xi_{0}^{2}$ system there is also a small magnetic field around $J\sim\Delta_{0}$ and $\alpha\sim 3/\xi_{0}$. While Eq. () 9 implies that for an antivortex pattern there is no circulating current and so no magnetic field is generated, we find numerically that the antivortices deviate a little from an ideal antivortex and there is sufficient circulating current to generate a magnetic field. For the $(5\times 5)\xi_{0}^{2}$ system there is almost no magnetic field. In general, most of the current density is concentrated close to the superconducting leads, so in the vortex phase $\alpha<1.25/\xi_{0}$ there is very little current and hence only a tiny magnetic field generated. When there are multiple singularities, $\alpha>1.25/\xi_{0}$, the currents around each singularity counterpropagate, resulting in the overall magnetic field vanishing (see Sec. III in the supplement [23]). For (e-g), the peak in $B_{\rm anom}$ around $J=\Delta_{0}$ is due to the fact that $f_{z}\sim J/(J^{2}+\omega^{2})$ (with $\omega$ the Matsubara frequency) [23]i.e., the magnetic field must peak at some finite $J$ and then decay for large $J$.
@@
Characterization and manipulation— Because circulating currents produce a uniform magnetic field, a natural approach to characterizing these vortices is via nano-magnetometry. Scanning probe techniques are well-established and can operate at cryogenic temperatures with high spatial resolution. Magnetic force microscopy has already been used to study conventional Josephson vortices [33, 34], and can operate with a resolution of 10 — 50 nm [35], or even atomic resolution [36]. Scanning nitrogen-vacancy center [37] and nanoSQUID methods [38] can now both achieve spatial resolutions of around 50 nm; indeed, recently, scanning nitrogen-vacancy centers were used to study vortices in NbSe2 [39]. These magnetometry techniques are complicated by the need to subtract the background ferromagnetic signal. One possible solution is to include a back gate to control the Rashba interaction [40]. One could then effectively switch off SOC to isolate the background ferromagnetic signal. The absence of SOC can, for example, be confirmed via the absence of a Josephson diode effect.
Such gate-based techniques for controlling Rashba SOC could also be used to manipulate these vortices. Consider a large weak link which has only weak (or absent) Rashba SOC. One can define an island within the weak link with strong Rashba SOC via a small back gate. We show in Fig.  4 examples where such Rashba “islands” can be used to control the size, position, and shape of a JAV. Such methods could be used to realize JAVs in arbitrarily sized junctions, overcoming the necessity of defining a square geometry of a coherence length. Indeed, several different vortices could be defined within a single weak link using this technique.
Outlook— We describe the characterization and manipulation of Josephson anomalous vortices, which appear in Josephson junctions where the weak link is a weak ferromagnet with strong spin-orbit coupling. The exchange field and spin-orbit coupling conspire to lead to circulating currents that we associate phenomenologically to a previously unexplored rotary term in the free energy. The same ingredients lead to textures in the triplet components ${\bm{f}}_{t}$.
@@ Acknowledgements.
I Appendix A: Lifshitz and rotary invariants in Ginzburg-Landau theory
Here we derive the rotary invariant in the Ginzburg–Landau picture phenomenologically, and discuss its meaning in more detail.
The expression for the free energy must remain unchanged with respect to various global transformations that leave the physical system invariant [20]. The transformations of the order parameter field are $\psi\mapsto\psi^{*}$ (time reversal $\mathcal{T}$, TR), $\psi({\bm{r}})\mapsto\psi(-{\bm{r}})$ (inversion $\mathcal{I}$), $\psi({\bm{r}})\mapsto\psi(R_{ij}r_{j})$ (global rotation), and $\psi({\bm{r}})\mapsto{}e^{i\varphi}\psi({\bm{r}})$ (global $U(1)$); coefficients transform keeping $\mathcal{F}$ invariant and real. Further restrictions would come from crystal symmetry, but we do not consider that here.
Expanding in gradients and expressing the result in terms of real coefficients, the possible terms with one gradient are
$\mathcal{F}_{1}$ $=d_{k}i[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]+c_{k}\partial_{k}|\psi|^{2}$ (11)
Then, $d_{k}$ is a TR-odd vector ($\mathcal{T},\mathcal{I}=-,$-) and $c_{k}$ a TR-even vector ($\mathcal{T},\mathcal{I}=+,$-). From integration by parts, the $c_{k}$ term is equivalent to the Ginzburg–Landau $a|\psi|^{2}$ term, and a similar boundary term $n_{k}c_{k}|\psi|^{2}$ where $n_{k}$ is the boundary normal vector.
The possible terms with two gradients are
$\mathcal{F}_{2}=$ $D_{jk}\partial_{j}\psi\partial_{k}\psi^{*}+ie_{jk}\partial_{j}\psi\partial_{k}\psi^{*}+f_{jk}\partial_{j}\partial_{k}|\psi|^{2}$
$+g_{jk}i\partial_{j}[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]$ (12)
where without loss of generality we have taken $D_{jk}=D_{kj}$, $e_{jk}=-e_{kj}$, and $f_{jk}=f_{kj}$, $g_{jk}=g_{kj}$. Then, $D_{jk}$ is a TR-even tensor ($\mathcal{T},\mathcal{I}=+,+$) and $e_{jk}$ a TR-odd tensor ($\mathcal{T},\mathcal{I}=-,+$). Similarly: $f_{jk}$ is a TR-even and $g_{jk}$ a TR-odd tensor.
For spatially constant coefficients, the terms aside from the superfluid stiffness $D_{jk}$ part are total derivatives. From integration by parts, the $f_{jk}$ term is equivalent to the $c_{k}$ term and a boundary term $n_{j}f_{jk}\partial_{k}|\psi|^{2}$. The other part $g_{jk}$ is equivalent to a Lifshitz invariant $(\partial_{j}g_{jk})i[\psi\partial_{k}\psi^{*}-\mathrm{c.c}]$ and a similar boundary term $n_{j}g_{jk}i[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]$. The contribution from the $e_{jk}$ term also reduces as
$\int_{M}ie_{jk}\partial_{j}\psi\partial_{k}\psi^{*}=$ $-\int_{\partial M}\frac{1}{2}n_{j}e_{jk}i[\psi\partial_{k}\psi^{*}-\psi^{*}\partial_{k}\psi]$ (13)
$+\int_{M}\frac{1}{2}(\partial_{j}e_{jk})i[\psi\partial_{k}\psi^{*}-\psi^{*}\partial_{k}\psi]\,.$ (14)
Consequently, also $e_{jk}$ is equivalent to a bulk Lifshitz invariant and a similar boundary term.
On a phenomenological level, the $e_{jk}$ term can then always as well be expressed in terms of a boundary term. However, as shown in Sec. I of the supplement [23], an equivalent term arises naturally as a bulk contribution in the microscopic theory, which is valid for arbitrary smooth spatially varying gauge and exchange fields (i.e., on scales longer than the mean free path). For interfaces described by such theory, the effective boundary Lifshitz invariant that is responsible for the vorticity is microscopically generated by the bulk $e_{jk}$ term.
II Appendix B: Forming a vortex due to the rotary invariant in a superconducting disk
The rotary invariant can lead to a formation of superconducting vortices. Consider the effect of a spatially constant rotary invariant $e_{xy}=-e_{yx}=\epsilon$ in a superconducting disk $M$ of radius $R$, in the absence of the Lifshitz invariant and for a spatially uniform and rotationally symmetric superfluid weight $D_{jk}=D_{S}\delta_{jk}$. The total free energy is
$F=$ $\int_{M}d^{2}r\left(-a|\psi|^{2}+b|\psi|^{4}+D_{S}|\nabla\psi|^{2}\right)$
$+\epsilon\int_{\partial M}\hat{n}\cdot{\rm Im}[\psi^{*}\nabla\times(\psi\hat{u}_{z})]$ (15)
$=$ $\int_{0}^{2\pi}d\varphi\int_{0}^{R}rdr\left(-a|\psi|^{2}+b|\psi|^{4}+D_{S}|\nabla\psi|^{2}\right)$
$+\epsilon\int_{0}^{2\pi}d\varphi{\rm Im}[\psi^{*}\partial_{\varphi}\psi],$ (16)
where the second line is written in polar coordinates. Note that we have chosen a convention where $a>0$ corresponds to the superconducting state.
The free energy of a spatially constant superconducting order parameter satisfying the Ginzburg-Landau equations, i.e., $\psi=\psi_{0}\equiv\sqrt{a/(2b)}$, is $F_{0}=-\pi R^{2}a^{2}/(4b)$. Let us check if the rotary invariant can provide an inhomogeneous state with a lower energy. We make the variational Ansatz
$\psi(r,\varphi)=\psi_{0}\tanh(r/\xi)e^{in\varphi},\quad n\in{\mathbb{Z}},$ (17)
ensuring that $\psi$ is uniquely defined at all positions inside the disk. Here $\xi$ is an Ansatz length scale optimized below. Up to exponentially small corrections for $R\gg\xi$, the change $\delta F=F-F_{0}$ reads
$\delta F=\frac{\pi a}{b}\left[a\xi^{2}c_{0}+D_{S}\left(c_{1}+n^{2}\log\frac{c_{2}R}{\xi}\right)+\epsilon n\right],$ (18)
where $c_{0}=(1+20\log 2)/12$, $c_{1}=(4\log 2-1)/3$ and $c_{2}\approx 1.25$. The term proportional to $a$ comes from the combination of the $a$ and $b$ terms in the free energy density, when using the particular value for $\psi_{0}$. Minimizing this with respect to $\xi$ yields $\xi=|n|\sqrt{D_{S}/(c_{0}a)}$ and for this minimum scale the free-energy change reads
$\delta F=\frac{\pi a}{b}\left\{D_{S}[c_{1}+n^{2}(1+\log\frac{c_{2}R}{\xi})]+\epsilon n\right\}.$ (19)
This variational Ansatz shows that the rotary invariant can drive the system to an inhomogeneous vortex state. The transition to the vortex state (i.e., with $\delta F<0$) for $\epsilon>0$ takes place with n=1n=-1, when
$\epsilon\geq D_{S}\left(1+c_{1}+\log\frac{c_{2}R}{\xi}\right).$ (20)
In other words, the vortex transition requires that the rotary invariant essentially exceeds the superfluid weight.
Note that this result is based on a crude variational Ansatz, so Eq. () 20 is an approximate estimate for the critical value of the rotary invariant.